OpenStudy (anonymous):

the area of a circle is 20.25pi cm^2. find the diameter.

Directrix (directrix):

The formula for the area of a circle is this: A = π · r² 20.25 π = π · r² Solve for r. Remember to double the value of r because twice the radius is the diameter. @Johnny_Reyes

OpenStudy (anonymous):

@Directrix does D=9? Thanks

OpenStudy (sleepyhead314):

yes D = 9 @Johnny_Reyes :)